OpenStudy (anonymous):

Which fraction is equivalent to 36%? A. 36//10 B. 9/25 C. 360/100 D. 3 6/10

OpenStudy (anonymous):

i think its either a or c

OpenStudy (texaschic101):

36% = 0.36 = 36/100 = 18/50 = 9/25

OpenStudy (anonymous):

ok thx i was confused

OpenStudy (texaschic101):

no problem...it sometimes can get confusing